PROGRESS TOWARDS HOLLOW WAVEGUIDE-ENHANCED LASER-INDUCED FLUORESCENCE

\underline{DANIEL B. HOLLAND} and GEOFFREY A. BLAKE, California Institute of Technology, Pasadena, California 91125.

Hollow waveguides (HWGs), essentially hollow fiber optics, have an open central core through which light and sample gas can travel coaxially. This enhanced light/sample interaction geometry has previously been utilized by others to perform laser-based absorption measurements within HWGs. The current work aims to achieve HWG-enhanced infrared laser-induced fluorescence (LIF), where both the laser beam and fluorescence would be efficiently waveguided though a tube of variable length. It is hoped that HWGs could be utilized to construct sensitive, small LIF spectrometers. An instrument employing HWG-LIF could potentially be fed sample material from a GC instrument or be used in a standalone configuration, for example. This talk will cover our progress towards this goal, with particular attention to LIF spectroscopy of water isotopologues (H_{2}O and HDO) using commercial telecom lasers.